Storrs Cycles has just started selling the new Cyclone mountain bike, with monthly sales as shown in the table. Month Sales JPlease answer part 2 of B
Answer #1

Moving average forecast = Average of previous three months

Forecast for May = (Feb + Mar + April) / 3

= (372 + 410 + 370) / 3 = 384

Do give a positive rating if you think this answer helped